"You purchased an airplane for $494,000 and will depreciate it using a 7-year MACRS with a 5-year life. Salvage value in year 5 is expected to be $188,000. The airplane is expected to increase revenues by $193,000 per year. However, O&M costs are expected to be $29,000 per year. Your company is in the 21% tax bracket and your MARR is 20%. What is the Net Present Worth of this investment?"
Answer #1

Using excel to calculate NPV

A B C D E F
Year 0 1 2 3 4 5
1 Airplane 494000
2 Increase in revenue 193000 193000 193000 193000 193000
3 O & M Costs 29000 29000 29000 29000 29000
4 Depreciation Macrs Rate 7 year 14.29% 24.49% 17.49% 12.49% 8.93%
5 Depreciation 70592.60 120980.60 86400.60 61700.60 44114.20
6 EBIT=Revenue-O&M Costs-Depreciation 93407.40 43019.40 77599.40 102299.40 119885.80
7 EAT=EBIT*(1-Tax Rate) 73791.846 33985.326 61303.526 80816.526 94709.782
8 Add Depreciation 70592.60 120980.60 86400.60 61700.60 44114.20
9 Add After tax Salvage Value 148520 (=188000*(1-21%)
10 Free Cash Flow -494000 144384.446 154965.926 147704.126 142517.126 287343.982
11 NPV $3,618.9485 Using excel=NPV(0.20,B10:F10)+A10

NPV =3619.95

Please discuss in case of doubt